-
#160
by
Space Pete
on 31 Jul, 2010 14:28
-
From ISS On-Orbit Status Report for 30/07/2010:In DC-1, FE-3 Mikhail Kornienko terminated the discharge process on the second 825M3 Orlan battery pack in the ZU-S recharge unit.
Other post EVA-25 closeout activities included:
• Mikhail & FE-5 Fyodor Yurchikhin configured the SM
PkhO (Transfer Compartment) and DC-1 to Nominal status.
• CDR Alexander Skvortsov prepared Russian tools to be used in the
upcoming US EVA-15 (ratchet wrench, extender, tether,
protective blanket with latch, etc.).
----------
Shannon Walker's 4th Blog Entry from the ISS: "Fixing the toilet. Or, what makes one appreciate simple mechanical systems and the local hardware store".
http://blogs.chron.com/inorbit/2010/07/fixing_the_toilet_or_what_makes_one_appreciate_simple_m_1.html----------
From ISS On-Orbit Status Report for 31/07/2010:In DC-1, FE-3 Mikhail Kornienko terminated the discharge process on the third 825M3 Orlan battery pack in the ZU-S recharge unit.
-
#161
by
rdale
on 01 Aug, 2010 00:19
-
Problems on ISS!
Sleep time was 3 hours ago so I found it interesting that there was chatter on the feed...
Loop Alpha ETCS pump module power was pulled because a short was detected by the RPC that provides power to it. Looking at data dump to see if it's a real problem with overcurrent. If it's not, they are fine and can power back up. It'll take about 30 minutes to get that data.
Still have good cooling for critical equipment in the lab. Node 2 & Node 3 have lost MTL, doing some ground powerdowns. MELFI being powered off.
-
#162
by
rdale
on 01 Aug, 2010 00:21
-
In the meantime shift 2 just came onboard, so they have extra hands on deck. While data is coming down they are making two plans:
1) Powerup methods in case this is just a glitch
2) Alternative power if loop alpha is really going offline
-
#163
by
rdale
on 01 Aug, 2010 00:42
-
Continuing to do powerdowns as data now completely on the ground for analysis.
IF they cannot get RPC closed, they will perform Loop A Shutdown from Warning handbook 2.674
-
#164
by
rdale
on 01 Aug, 2010 01:13
-
I'm not in a position to listen any more tonight, if anyone else wants to help out - be my guest! Unfortunately the archive system is down so nothing will be saved, and the system doesn't accept WAV files.
-
#165
by
rdale
on 01 Aug, 2010 01:16
-
They started sending an update just as I typed that...
Loop A ETCS pump is the problem. The RPC is what tripped. We saw evidence that it is the pump itself, and that's bad news. Not for sure, but saw indications.
-
#166
by
rdale
on 01 Aug, 2010 01:16
-
RPC tripped to give us protection, which is why there are so many error messages. We might reactivate the RPC to see if the pump will restart, but won't do that unless we are 100% sure that it can't make things worse.
-
#167
by
rdale
on 01 Aug, 2010 01:17
-
We closed the comm gaps so we have 4 hours of continuous comm. We may have you move stuff from MELFI-1 to MELFI-2 to save some science, but that would be the only other needed crew interaction.
-
#168
by
Targeteer
on 01 Aug, 2010 01:57
-
On comms still--they are really past their bedtime. (Power?) Loop A is being shut down and they mentioned going to two CMGs thus T2 is being isolated. Also discussing a get to sleep plan.
-
#169
by
Targeteer
on 01 Aug, 2010 02:54
-
JEM caution warning system has been deactivated and MELFI 1 to 2 transfer is required and being passed up by voice. Nutrition samples are being transferred. MELFI 1 is being emptied. Node 2 MDM is about to go down--no cautions expected.
-
#170
by
rdale
on 01 Aug, 2010 03:40
-
Still working on moving all samples from MELFI-1 to MELFI-2. A little confusion on what samples are in what drawers, but seems like things are flowing. They put them in cold bags apparently using "ice bricks" and goal is to take anything out of the sections.
-
#171
by
rdale
on 01 Aug, 2010 03:44
-
-
#172
by
rdale
on 01 Aug, 2010 03:46
-
MELFI transfers are complete. Still marching down the procedure to shut things down. Ground is looking at tomorrow's timeline to shift things that they can so crew can sleep as long as possible.
One thing was on the schedule for 8am, if the crew does it now (just a 5 minute procedure) then they won't need to wake them up. Should be able to get to bed in about 5-10 minutes.
-
#173
by
rdale
on 01 Aug, 2010 03:52
-
Few more small issues - In Columbus, power down PWS1 and SSC7. In JEM, make sure window shutters are closed.
-
#174
by
Chris Bergin
on 01 Aug, 2010 03:56
-
Good work with the updates Rob and Targeteer!
-
#175
by
rdale
on 01 Aug, 2010 04:06
-
Working in Node 1 Port 4 rack area, they will be performing procedure from ISS IFM 2.3.405 (don't recall ever seeing the manual.) Need to get a jumper and temp stow it there for later work.
-
#176
by
rdale
on 01 Aug, 2010 04:08
-
Ground is not final on doing that tonight, leaning towards tomorrow morning, looking for crew input for how long that takes.
Ground is about to start spinning down CMG 1 & 4.
[Jumping way ahead of the curve, but I do recall a spare pump module was sent up a few flights back?]
-
#177
by
Targeteer
on 01 Aug, 2010 04:08
-
CMGs 1 and 4 are about to spin down...
-
#178
by
rdale
on 01 Aug, 2010 04:14
-
Thanks for the assist Targeteer - it's too late for me already so I'll let you finish up if they have more words tonight...
I've attached a zipfile of some of the callup audio.
-
#179
by
The-Hammer
on 01 Aug, 2010 04:34
-
Jumping way ahead of the curve, but I do recall a spare pump module was sent up a few flights back?
There are four spare pump modules on-orbit:
1 each on ESP-2 (STS-121), ESP-3 (STS-127), ELC-1 (STS-129), ELC-2 (STS-129).